OpenStudy (anonymous):

2x^2tan(x)/sec(x) find f'(x) and f'(3)

zepdrix (zepdrix):

Are the tangent and secant both in the exponent? :O

zepdrix (zepdrix):

\[\frac{ 2x^2tanx }{ secx }\] Errr I suppose it's like this? :D that would make more sense.

OpenStudy (anonymous):

yeah thats it, for some reason my open study isnt working right, so i hope it works because I really need the help

zepdrix (zepdrix):

If we try to differentiate it from here, it'll be a bit of a mess, because we have the product rule within the quotient rule... Rewrite tangent and secant in terms of sines and cosines, and you should see a nice cancellation. :D Remember the identities?

OpenStudy (anonymous):

sec=1/cos and tan=sin/cos?

zepdrix (zepdrix):

mhm c:

zepdrix (zepdrix):

So you're dividing by secant, when you change it, using that identity, you end up dividing by a fraction. Soooo what can we do with that? :D Remember how to flip it? XD

OpenStudy (anonymous):

haha yup multiply it by cos/1

zepdrix (zepdrix):

Anything cancel out? :3

OpenStudy (anonymous):

the two cos?

zepdrix (zepdrix):

yay

zepdrix (zepdrix):

\[\frac{ 2x^2tanx }{ secx }=2x^2sinx\] Wow so we were able to simplify it quite nicely :O So from here you want to give it a try or still confused? :) Product rule stuffs.

OpenStudy (anonymous):

im gonna give her a go.

OpenStudy (anonymous):

if we use the product rule (4x^1) and the derivative of sinx is cosx

zepdrix (zepdrix):

Mhm sounds right, come up with an answer? :D

OpenStudy (anonymous):

so would f'(x)= (4x)(cos(x))

zepdrix (zepdrix):

Woops D: You just differentiated them both, uh ohhhhh. We have the product of 2 functions of x. (2 things with x). So we have to apply the product rule. \[(uv)'=u'v+uv'\] \[[2x^2tanx]'=(2x^2)'tanx+2x^2(tanx)'\] Understand what you need to do? The primes signify a place where you still need to take a derivative.

OpenStudy (anonymous):

(4x)(tan(x))+2x^2(sec(x))^2 is that what you mean? sec(x)^2 = derivative of tan(x)

zepdrix (zepdrix):

Yes good job! c:

OpenStudy (anonymous):

?? is that the answer

zepdrix (zepdrix):

Yes.

OpenStudy (anonymous):

That answer didnt work in my online assignment. I put it in www.wolframalpha.com and it says f ' (x) = 2 x (x cos(x)+2 sin(x))